SUB-DOPPLER SPECTROSCOPY OF trans-HOCO RADICAL IN THE OH STRETCHING MODE

\underline{CHIH-HSUAN CHANG}, GRANT BUCKINGHAM, and DAVID J. NESBITT, JILA, National Institute of Standards and Technology University of Colorado, and Department of Chemistry and Biochemistry University of Colorado at Boulder, Colorado 80309.

Rovibrational spectroscopy of the fundamental OH stretching mode of trans-HOCO radical has been studied via sub-Doppler high resolution infrared laser absorption in a discharge slit-jet expansion. The trans-HOCO radical is formed by discharge dissociation of H_{2}O to form OH, which then combines with CO and cools in the Ne expansion to a rotational temperature of 16 K. Rigorous assignment of both a-type and b-type spectral transitions is made possible by 2-line combination differences from microwave studies^{a}, with full rovibrational analysis of the spectrum based on a Watson asymmetric top Hamiltonian. Additionally, fine structure splittings of each line due to electron spin are completely resolved and thus permitting all three ϵ_{a}_{a}, ϵ_{b}_{b}, ϵ_{c}_{c} spin rotation constants to be experimentally determined. Furthermore, the ratio of transition dipole moment along the molecule A and B axis is determined to be μ_{a}/μ_{b} = 1.72(5), in agreement with quantum theoretical calculations(1.85).

aT. Oyama,W Funato, Y. Sumiyoshi,and Y. Endo, J. Chem. Phys. 134, 174303.